A soup can is in the shape of a cylinder. The can has a volume of 342 cm^3 and a diameter of 6cm.
(A)Express the height of the can in terms of pi. [Show all work]

(B)Determine the maximum number of soup cans that can be stacked on their base between two shelves if the distance between the shelves is exactly 36 cm. [Explain your answer]

Answer by lwsshak3(11628)

A soup can is in the shape of a cylinder. The can has a volume of 342 cm^3 and a diameter of 6cm.
(A)Express the height of the can in terms of pi. [Show all work]
volume=h*area of base
height=volume/area of base
=342/(πd^2/4)
=342/(π6^2/4)
=342/9π
height=38/π cm
(B)Determine the maximum number of soup cans that can be stacked on their base between two shelves if the distance between the shelves is exactly 36 cm. [Explain your answer]
height=38/π cm
3*height=3*38/π=36.287.. cm
Therefore, maximum number of soup cans that can be stacked on their base=2